# Motion-compensated cardiac MRI using low-rank diffeomorphic flow (DMoCo)

### v2

@@ Abstract
We introduce an unsupervised motion-compensated image reconstruction algorithm for free-breathing and ungated 3D cardiac magnetic resonance imaging (MRI). We express the image volume corresponding to each specific motion phase as the deformation of a single static image template. The main contribution of the work is the low-rank model for the compact joint representation of the family of diffeomorphisms, parameterized by the motion phases. The diffeomorphism at a specific motion phase is obtained by integrating a parametric velocity field along a path connecting the reference template phase to the motion phase. The velocity field at different phases is represented using a low-rank model. The static template and the low-rank motion model parameters are learned directly from the k-space data in an unsupervised fashion. The more constrained motion model is observed to offer improved recovery compared to current motion-resolved and motion-compensated algorithms for free-breathing 3D cine MRI.
@@ I Introduction
Magnetic Resonance (MR) is the gold standard for evaluating cardiac function. The 2D cardiac cine acquisition strategy requires extensive preparatory scans for slice localization and requires breath hold, which many subject groups (e.g., children or patients with reduced tidal volume) cannot comply with. 3D free-breathing and ungated acquisition schemes that offer isotropic spatial resolution and eliminate the need for preparatory scans were recently introduced [17, 23, 9, 29, 12, 18, 10, 26, 15, 27, 2]. These schemes estimate the respiratory and cardiac phases from k-space center, dedicated k-space navigators, or ECG signals. The k-space data is grouped into cardiac and respiratory bins using the above phase information, followed by the joint reconstruction of the images using total variation or low-rank penalties. Although these motion-resolved approaches offer great promise, the main challenge is the associated long acquisition times to ensure good sampling of the bins. Although supervised deep learning (DL) approaches may offer improved trade-offs, the challenge is the lack of training data; unlike 2D settings, the acquisition of fully sampled images is not possible in our 5D setting. In addition, the extension of current DL methods is challenging because of the enormous memory demand. Recently, patient-specific self-supervised deep learning methods were introduced [30, 32, 28, 13, 16, 33] to further reduce data demand in motion-resolved 4D (2D+cardiac and respiratory dimensions) MRI. These methods express the image at a specific cardiac and respiratory phase as a nonlinear function of the low-dimensional signals (e.g. cardiac and respiratory motion). In our 5D setting, a large 3D CNN is needed to represent the image volumes in each phase; training of this model using backpropagation will be associated with high memory demand.
@@
Motion-compensated (MoCo) reconstruction schemes that explicitly model the deformations between the phases are significantly more data efficient than the above motion resolved strategies [21, 19, 24, 31]. Because they combine the information from the different bins, they are less sensitive to varying numbers of samples in each bin. The main drawback of these methods is the high computational complexity of image registration. Deep learning (DL) methods to estimate respiratory motion [19] and unrolled DL methods that incorporate motion compensation have been recently introduced [22]. These methods are not directly applicable to our 5D setting due to the extreme undersampling of the bins, the high memory demand, and the need for fully-sampled reference data. An unsupervised DL approach termed MoCo-SToRM [34] was recently introduced for respiratory motion-compensated lung MRI. MoCo-SToRM models the volume in each respiratory phase as the deformed version of a static image template. The motion field at each phase is modeled as the output of a CNN in response to low-dimensional latent vectors that correspond to the phase. The latent vectors, the CNN parameters, and the image template are jointly estimated from the undersampled k-t space data of each subject, similar to self-supervised motion-resolved strategies[30, 32, 28, 13, 16, 33].
We introduce an unsupervised approach for the recovery of motion-compensated cardiac cine acquired using free breathing and ungated 3D-SSFP acquisitions. Similar to MoCo-SToRM [14], we represent the image in each phase (cardiac and respiratory) as the deformed version of a static image template. The main distinction of the proposed scheme is the more constrained low-rank diffeomorphic motion model. The deformation model in MoCo-SToRM does not guarantee that the deformation of the heart and surrounding tissues maintains integrity and connectivity, preventing any intersections between different regions. The lack of regularity constraints translates to inaccurate motion estimates when the phases are heavily undersampled (e.g, around 100x or more). We are motivated by diffeomorphic models used in medical image registration that represents the deformation between two images as an invertible function, such that both the function and its inverse are continuously differentiable [3, 8], i.e., smooth. These schemes express the deformation as the time integral of a time-varying vector field, which guarantees the above properties.
@@
The main contribution of this work is the low-rank model for the compact joint representation of the family of diffeomorphisms. This is a generalization of current diffeomorphic models, which consider the registration of a pair of images, to the joint registration of a family of images. In particular, we model the deformation at each motion phase as the integral of a phase-dependent velocity field along a linear path connecting the template phase and the desired phase. We note that the collection of the velocity fields at different motion phases is significantly more low-rank than the original images and even the deformation fields at different phases; the low-rank velocity model thus offers a compact representation of the diffeomorphisms and hence the images. We additionally incorporate regularization prior to encourage the diffeomorphisms to be independent of the integration path, which further reduces the degree of freedom. We estimate the velocity fields and the static template from the measured k-space data in an unsupervised fashion. The coupling of the motion fields at different phases via the overlapping path integrals, the use of velocity constraints to make the path of the deformations independent, and the low-rank model make the deformation model more constrained than the unsupervised approach in [34] and the supervised approach [19]. This enables us to extend the MoCo approach to the cardiac and respiratory resolved recovery from six minutes of acquisition. This work builds on our previous conference paper [14], which used the MoCo-SToRM [34] motion model for 5D cardiac MRI.
@@ II-A Multichannel MRI Acquisition
Here, $\mathbf{r}=(x,y,z)\in\mathbb{R}^{3}$ denotes the spatial variable and $t$ denotes the time at which the k-space measurements $\mathbf{b}(t)$ were acquired. $\rho_{t}$ denotes the spatial volume at time instant $t$ and $\mathcal{A}_{\mathbf{k}(t)}$ denotes the multichannel Fourier transform corresponding to the k-space locations $\mathbf{k}(t)$. $\mathbf{n}(t)$ denotes the measurement noise of variance $\sigma^{2}$.
@@ II-B Diffeomorphic Image Registration
Given two $n$-dimensional image volumes $I_{0},I_{1}\in\mathbb{R}^{n}\rightarrow\mathbb{C}$, image registration aims to find a deformation $\phi_{1}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ such that $I_{1}(\mathbf{r})=I_{0}(\phi_{1}(\mathbf{r}))$. A deformation may be evaluated as $\phi_{1}(\mathbf{r})=\mathbf{r}+\boldsymbol{u}(\mathbf{r})$, where $\boldsymbol{u}$ denotes the displacement. Although its inverse may be approximated as $\mathbf{r}=\phi_{1}(\mathbf{r})-\boldsymbol{u}(\phi_{1}(\mathbf{r}))$, this is only as exact when $\boldsymbol{u}(\mathbf{r})\rightarrow\boldsymbol{u}(\phi(\mathbf{r}))$; that is, for small deformations that satisfy $\boldsymbol{u}\rightarrow 0$. One can compose several infinitesimal deformations
$\phi_{N}(\mathbf{r})=(\mathbf{r}+\boldsymbol{u}_{N}(\mathbf{r}))\leavevmode\nobreak\ \circ\leavevmode\nobreak\ \ldots\leavevmode\nobreak\ \circ\leavevmode\nobreak\ (\mathbf{r}+\boldsymbol{u}_{0}(\mathbf{r})),$
to obtain a large deformation that is invertible. This intuition has led to the representation of diffeomorphisms, which are often expressed as an endpoint of a flow. $\phi_{\delta}(\mathbf{r}):[0,1]\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, associated with a smooth $n$-dimensional vector field $\boldsymbol{\nu}_{\delta}(\mathbf{r}):[0,1]\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, specified by the ordinary differential equation [6, 8]
@@ III Low-rank diffeomorphic Image Model
Unlike motion-resolved methods that group the readouts into discrete bins and jointly recover the binned images, we represent the image volume $\rho_{t}(\mathbf{r})$ at the time instant $t$ as the function of the continuous phase variable $\boldsymbol{\tau}(t)\in\mathbb{R}^{2}$.
@@
III-A Parametric Diffeomorphic Motion Model
Most diffeomorphic registration methods consider the registration of a pair of images. In this work, we consider the joint registration of images $\rho_{t}(\mathbf{r})$ using a parametric family of diffeomorphisms, parameterized by $\boldsymbol{\tau}$. The diffeomorphisms are represented in terms of a parametric velocity tensor $\mathbf{v}_{\boldsymbol{\tau}}\Big{(}\frac{d\boldsymbol{\tau}}{dt},\mathbf{r}\Big{)}$ at the motion phase $\boldsymbol{\tau}$. The tensor $\mathbf{v}_{\boldsymbol{\tau}}$ can be thought of as the concatenation of two velocity fields. For instance, the first component of the tensor denoted by $\mathbf{v}_{\boldsymbol{\tau}}(0,\mathbf{r})$ is the tissue velocity at the phase $\boldsymbol{\tau}$, when the cardiac phase increases at the rate $dc/dt=1$. When both phases are increased at a speed $\boldsymbol{\tau}^{\prime}=\begin{bmatrix}dc/dt&dr/dt\end{bmatrix}^{T}$, we assume linearity and obtain the tissue velocity as
@@ III-A Parametric Diffeomorphic Motion Model
We now consider a curve $\boldsymbol{\gamma}:[0,1]\rightarrow\mathcal{C}_{\boldsymbol{\tau}}$ in phase space (see Fig. 1) starting at $\mathbf{0}$ and ending at $\boldsymbol{\tau}$ with $\boldsymbol{\gamma}(0)=\mathbf{0}$ and $\boldsymbol{\gamma}(1)=\boldsymbol{\tau}$ to define the diffeomorphism $\varphi_{\boldsymbol{\tau}}(\mathbf{r})$. Here, $\boldsymbol{\gamma}(\delta)$ are points on the curve and $\delta$ is an arbitrary parameterization. For the curve $\mathcal{C}_{\boldsymbol{\tau}}$, we consider the one-parameter velocity field:
@@
The above formulation produces parametric deformations $\varphi_{\boldsymbol{\tau}}$ that are dependent on the specific curve $\mathcal{C}_{\boldsymbol{\tau}}$. In this work, we define $\mathcal{C}_{\boldsymbol{\tau}}$ as a straight line-segment, connecting $\mathbf{0}$ and $\boldsymbol{\tau}$. As discussed in Section II-B, each of the deformation fields $\varphi_{\boldsymbol{\tau}}$ can be inverted by running the ODE backwards as in (3). Thus, the image volume at phase $\tau_{1}$ can be deformed to phase $\tau_{2}$ using the transformation $\varphi_{\boldsymbol{\tau}_{2}}\circ\leavevmode\nobreak\ \varphi_{\boldsymbol{\tau}_{1}}^{-1}$, where $\circ$ is the composition mapping.
III-B Path constraint on velocity fields
@@
With no additional constraints on the velocity field, we may have $\psi_{\boldsymbol{\tau}}\neq\varphi_{\boldsymbol{\tau}}$. If the integrals were path independent, a diffeomorphism could be directly evaluated from $\boldsymbol{\tau}_{1}$ to $\boldsymbol{\tau}_{2}$ by integrating the velocity field on a path connecting the two points in phase space rather than evaluating it as $\varphi_{\boldsymbol{\tau}_{2}}\circ\varphi_{\boldsymbol{\tau}_{1}}^{-1}$.
We note that the constraint $\psi_{\tau}=\varphi_{\tau}$ is not necessary for the representation of the deformations and to realize the algorithm. However, in the interest of reducing the degrees of freedom in the representation, we propose to add a penalty to encourage the diffeomorphisms to be path independent. We consider a penalty involving the path-dependent expected difference in the deformations $\mathbb{E}_{\mathcal{D}_{\tau}}\|\psi_{\tau}-\varphi_{\tau}\|^{2}$. In practice, at each epoch we choose $\mathcal{D}_{\boldsymbol{\tau}}$ as a random perturbation of the line segment $\mathcal{C}_{\boldsymbol{\tau}}$. We randomize the perturbations over epochs to ensure that the relation holds for all paths. Our in silico results in Fig. 3 show the benefit of adding this constraint.
III-C Low-rank representation of velocity tensors
The deformations $\varphi_{\boldsymbol{\tau}}(\mathbf{r})$ are computed in terms of the velocity fields $\mathbf{v}_{\boldsymbol{\tau}}(\mathbf{r})$, parametrized by phase $\boldsymbol{\tau}$. There are extensive similarities between the parametric tissue velocities $\mathbf{v}_{\boldsymbol{\tau}}$ at different phases. For example, the deformations in different cardiac phases may be modeled by a relatively constant velocity field, as shown in Fig. 2. In the interest of realizing a compact representation, we propose to use a low-rank model to represent the velocity field:

### v3

# Motion-compensated recovery using diffeomorphic flow (DMoCo) for 5D cardiac MRI

@@ Abstract
We introduce an unsupervised motion-compensated image reconstruction algorithm for free-breathing and ungated 5D functional cardiac magnetic resonance imaging (MRI). The proposed algorithm models the 3D volumes at different cardiac and respiratory phase as deformed versions of a single static image template. The template as well as the parameters of the deformations are jointly estimated from the measured k-t space data. The main contribution of this work is a constrained representation of the deformations as a family of diffeomorphisms, parameterized by the motion phases. The deformation at a specific motion phase is derived as the integral of a parametric velocity tensor along a path connecting the reference template phase to the motion phase. The velocity tensor at different phases is represented using a low-rank model. The results show that the more constrained motion model offers improved recovery compared to current motion-resolved and motion-compensated algorithms for free-breathing and ungated 3D cardiac MRI data.
@@ I Introduction
Magnetic Resonance (MR) is the gold standard for evaluating cardiac function. Current clinical 2D cine acquisition strategy requires extensive preparatory scans for slice localization and requires breath hold, which many subject groups (e.g., children or patients with reduced tidal volume) cannot comply with. 3D free-breathing and ungated acquisition schemes that offer isotropic spatial resolution and eliminate the need for preparatory scans were recently introduced [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. These schemes estimate the respiratory and cardiac phases from the center of k space, dedicated k space navigators, or ECG signals. The k-space data are grouped into cardiac and respiratory bins using the above phase information, followed by joint reconstruction of the images using total variation or low-rank penalties. Although these motion-resolved approaches offer great promise, the main challenge is the associated long acquisition times to ensure good and even sampling of the bins. Although supervised deep learning (DL) approaches may offer improved trade-offs, the challenge is the lack of training data; unlike 2D settings, the acquisition of fully sampled images is not possible in our 5D setting. In addition, the extension of current DL methods is challenging because of the high memory demand in the 3D setting. Recently, patient-specific self-supervised deep learning methods were introduced [12, 13, 14, 15, 16, 17] to reduce data demand in motion-resolved 4D MRI (two spatial dimensions along with cardiac and respiratory). These methods express the image at a specific cardiac and respiratory phase as a nonlinear function of the low-dimensional signals (e.g. cardiac and respiratory motion). In our 5D setting, a large 3D CNN is needed to represent the image volumes in each phase; training of this model using backpropagation will be associated with high memory demand.
@@
Motion-compensated (MoCo) reconstruction schemes that explicitly model the deformations between the phases are significantly more data efficient than the above motion resolved strategies [18, 19, 20, 21]. Because they combine the information from the different motion phases, they are less sensitive to varying numbers of samples in each bin. The main drawback of these methods is the high computational complexity of image registration. Deep learning (DL) methods for estimating respiratory motion [19] and unrolled DL methods that incorporate motion compensation have recently been introduced [22]. These methods are challenging to apply directly to our 5D setting because of the extreme and uneven undersampling of the bins, the high memory demand, and the need for fully sampled reference data. An unsupervised DL approach called MoCo-SToRM [23] was recently introduced for respiratory motion-compensated lung MRI. MoCo-SToRM models the volume in each respiratory phase as the deformed version of a static image template. The motion field at each phase is modeled as the output of a CNN in response to low-dimensional latent vectors that correspond to the phase. The latent vectors, the CNN parameters, and the image template are jointly estimated from the undersampled k-t space data of each subject, similar to self-supervised motion-resolved strategies[12, 13, 14, 15, 16, 17, 24]. We note that MoCo-SToRM was originally developed for respiratory-compensated 4D MRI. The direct extension of the simple deformation model in MoCo-SToRM to our respiratory and cardiac motion-compensated setting can result in inaccurate deformation estimates, especially when the phases are heavily undersampled and the contrast-to-noise ratio is low. Although approaches similar to MoCo-SToRM [25, 26] have recently been introduced for the 5D setting, they often require ferumoxytol infusion that offers an order of magnitude improvement in contrast to the noise ratio to obtain reliable results [25]. Because ferumoxytol administration is not a viable option in many clinical settings, the development of methods for a low SNR and a highly accelerated setting is desirable.
@@
We introduce an unsupervised approach for the recovery of motion-compensated cardiac cine acquired using free breathing and ungated 3D-SSFP acquisitions. Similar to MoCo-SToRM [26], we represent the image in each phase (cardiac and respiratory) as the deformed version of a static image template. The main distinction of the proposed scheme with MoCo-SToRM is the more constrained low-rank diffeomorphic motion model. The deformation model in MoCo-SToRM does not guarantee that the deformation of the heart and surrounding tissues maintains integrity and connectivity and is devoid of intersections between different regions. We are motivated by diffeomorphic models used in medical image registration that represent the deformation between two images as an invertible function, so that both the function and its inverse are continuously differentiable [27, 28]. These schemes express the deformation as the time integral of a static vector field, which guarantees the above properties. The main contribution of this work is a low-rank model for the compact joint representation of the family of diffeomorphisms. This work involves the generalization of current diffeomorphic models [28], which consider the registration of a pair of images, to the joint registration of a family of images. In particular, we model the deformation at each motion phase as the integral of a phase-dependent velocity tensor along a linear path connecting the template phase and the desired phase. We observe that the collection of the velocity tensors at different motion phases is significantly more low-rank than the original images and the deformation fields at different phases. The low-rank velocity representation thus offers a compact model of the diffeomorphisms, and hence the images. We additionally incorporate a regularization prior to encourage the diffeomorphisms to be independent of the integration path, which further reduces the degree of freedom. We estimate the velocity tensors and the static template from the measured k-space data in an unsupervised fashion. The coupling of the motion fields at different phases via the overlapping path integrals, the use of velocity constraints the deformations independent of the integration path, and the low-rank model make the deformation model more constrained than the unsupervised approach in [23] and the supervised approach [19]. The preliminary studies in the paper shows the feasibility of DMoCo to enable cardiac and respiratory resolved recovery of 5D cardiac MRI data from six minutes of radial free-breathing and ungated acquisition. The main objective of this work is on introduce the DMoCo framework and to show its feasibility in this setting; future work would focus on a more thorough clinical validation.
@@ II-A Multichannel MRI Acquisition
Here, $\mathbf{r}=(x,y,z)\in\mathbb{R}^{3}$ denotes the spatial variable and $t$ denotes the time at which the k-space measurements $\mathbf{b}(t)$ were acquired. $\rho_{t}$ denotes the spatial volume at the time-instant $t$ and $\mathcal{A}_{\mathbf{k}(t)}$ denotes the multichannel Fourier transform corresponding to the k-space locations $\mathbf{k}(t)$. $\mathbf{n}(t)$ denotes the measurement noise of variance $\sigma^{2}$.
@@ II-B Diffeomorphic Image Registration
Given two $n$-dimensional image volumes $I_{0},I_{1}\in\mathbb{R}^{n}\rightarrow\mathbb{C}$, image registration aims to find a deformation $\phi_{1}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ such that $I_{1}(\mathbf{r})=I_{0}(\phi_{1}(\mathbf{r}))$. A deformation may be evaluated as $\phi_{1}(\mathbf{r})=\mathbf{r}+\boldsymbol{u}(\mathbf{r})$, where $\boldsymbol{u}$ denotes the displacement. Although its inverse may be approximated as $\mathbf{r}=\phi_{1}(\mathbf{r})-\boldsymbol{v}(\phi_{1}(\mathbf{r}))$, this is only as exact when $\boldsymbol{u}(\mathbf{r})\rightarrow\boldsymbol{v}(\phi(\mathbf{r}))$; that is, for infinitesimal deformations that satisfy $\boldsymbol{v}\rightarrow 0$. One can compose several infinitesimal deformations
$\phi_{N}(\mathbf{r})=(\mathbf{r}+\boldsymbol{v}_{N}(\mathbf{r}))\leavevmode\nobreak\ \circ\leavevmode\nobreak\ \ldots\leavevmode\nobreak\ \circ\leavevmode\nobreak\ (\mathbf{r}+\boldsymbol{v}_{0}(\mathbf{r})),$
to obtain a large deformation that is invertible. This intuition has led to the representation of diffeomorphisms, which are often expressed as the endpoint of a flow. $\phi_{\delta}(\mathbf{r}):[0,1]\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, associated with a smooth $n$-dimensional vector field $\boldsymbol{\nu}_{\delta}(\mathbf{r}):[0,1]\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, specified by the ordinary differential equation [29, 28]
@@ III Low-rank diffeomorphic Image Model
The diffeomorphic flow-based motion compensated (DMoCo) approach represents each volume in the time series as a deformed version of a static template. We now describe the main aspects of the framework, which are (a) image representation, (b) diffeomorphic motion model using a parametric velocity representation, and (c) low-rank model for the velocities.
III-A Image representation
We represent the image volume $\rho_{t}(\mathbf{r})$ at the time instant $t$ as a function of the continuous phase variable $\boldsymbol{\tau}(t)\in\mathbb{R}^{2}$.
@@ III-A Image representation
III-B Parametric Diffeomorphic Motion Model
The main distinction of DMoCo from MoCo-SToRM is the more constrained model used to represent $\varphi_{\boldsymbol{\tau}}(\mathbf{r})$. Most diffeomorphic registration methods consider the registration of a pair of images. In this work, we consider the joint registration of images $\rho_{t}(\mathbf{r})$ using a parametric family of diffeomorphisms, parameterized by $\boldsymbol{\tau}$. The diffeomorphisms are represented in terms of a parametric velocity tensor $\mathbf{v}_{\boldsymbol{\tau}}\Big{(}\frac{d\boldsymbol{\tau}}{dt},\mathbf{r}\Big{)}$ at the motion phase $\boldsymbol{\tau}$. The tensor $\mathbf{v}_{\boldsymbol{\tau}}$ can be thought of as the concatenation of two velocity fields. For example, the first component of the tensor denoted by $\mathbf{v}_{\boldsymbol{\tau}}(0,\mathbf{r})$ is the tissue velocity in the phase $\boldsymbol{\tau}$, when the cardiac phase increases at the rate $dc/dt=1$. When both phases are increased at a speed $\boldsymbol{\tau}^{\prime}=\begin{bmatrix}dc/dt&dr/dt\end{bmatrix}^{T}$, we assume linearity and obtain the tissue velocity as
@@ III-B Parametric Diffeomorphic Motion Model
We now consider a curve $\boldsymbol{\gamma}:[0,1]\rightarrow\mathcal{C}_{\boldsymbol{\tau}}$ in phase space (see Fig. 1) starting at $\mathbf{0}$ and ending at $\boldsymbol{\tau}$ with $\boldsymbol{\gamma}(0)=\mathbf{0}$ and $\boldsymbol{\gamma}(1)=\boldsymbol{\tau}$ to define the diffeomorphism $\varphi_{\boldsymbol{\tau}}(\mathbf{r})$. $\boldsymbol{\gamma}(\delta)$ are points on the curve, and $\delta$ is an arbitrary parameterization. For the curve $\mathcal{C}_{\boldsymbol{\tau}}$, we consider the one-parameter velocity tensor:
@@
The above formulation produces parametric deformations $\varphi_{\boldsymbol{\tau}}$ that are dependent on the specific curve $\mathcal{C}_{\boldsymbol{\tau}}$. In this work, we define $\mathcal{C}_{\boldsymbol{\tau}}$ as a straight line segment that connects $\mathbf{0}$ and $\boldsymbol{\tau}$. As discussed in Section II-B, each of the deformation fields $\varphi_{\boldsymbol{\tau}}$ can be inverted by running the ODE backwards as in (3). Thus, the image volume at phase $\tau_{1}$ can be deformed to phase $\tau_{2}$ using the transformation $\varphi_{\boldsymbol{\tau}_{2}}\circ\leavevmode\nobreak\ \varphi_{\boldsymbol{\tau}_{1}}^{-1}$, where $\circ$ is the composition mapping.
III-C Path constraint on velocity tensors
@@
With no additional constraints on the velocity tensor, we may have $\psi_{\boldsymbol{\tau}}\neq\varphi_{\boldsymbol{\tau}}$. If the integrals were path independent, a diffeomorphism could be directly evaluated from $\boldsymbol{\tau}_{1}$ to $\boldsymbol{\tau}_{2}$ by integrating the velocity tensor on a path connecting the two points in phase space rather than evaluating it as $\varphi_{\boldsymbol{\tau}_{2}}\circ\varphi_{\boldsymbol{\tau}_{1}}^{-1}$.
We note that the constraint $\psi_{\tau}=\varphi_{\tau}$ is not necessary for the representation of the deformations and to realize the algorithm. However, in the interest of constraining the representation, we propose to add a penalty to encourage the diffeomorphisms to be path independent. We consider a penalty involving the path-dependent expected difference in the deformations $\mathbb{E}_{\mathcal{D}_{\tau}}\|\psi_{\tau}-\varphi_{\tau}\|^{2}$. In practice, at each epoch we choose $\mathcal{D}_{\boldsymbol{\tau}}$ as a random perturbation of the line segment $\mathcal{C}_{\boldsymbol{\tau}}$. We randomize the perturbations over epochs to ensure that the relation holds for all paths. Our in silico results in Fig. 3 show the benefit of adding this constraint.
III-D Low-rank representation of velocity tensors
The deformations $\varphi_{\boldsymbol{\tau}}(\mathbf{r})$ are computed in terms of the velocity tensors $\mathbf{v}_{\boldsymbol{\tau}}(\mathbf{r})$, parametrized by phase $\boldsymbol{\tau}$. There are extensive similarities between parametric tissue velocities $\mathbf{v}_{\boldsymbol{\tau}}$ in different phases. For example, deformations in different cardiac phases (for a specific respiratory phase) may be approximated by a relatively constant velocity tensor, as shown in Fig. 2. In the interest of realizing a compact representation, we propose to use a low-rank model to represent the velocity tensor at a specific phase $\boldsymbol{\tau}$: